OpenStudy (mony01):

Solve the IVP. y'=3t^(2)-y y(0)=1 [0,1]

OpenStudy (freckles):

The is a linear diff equation y'+y=3t^2 multiply by v on both sides vy'+vy=v3t^2 where v is the integrating factor

OpenStudy (freckles):

you want to choose v such that you have the left hand side is (vy)'=vy'+v'y. So what does v' need to be in order for vy'+vy to be the same?

OpenStudy (mony01):

I dont understand what you're trying to say

myininaya (myininaya):

freckles is saying it is a linear differential equation

myininaya (myininaya):

and he/she is trying to get you to remember the product rule

myininaya (myininaya):

where v and y are functions of x (vy)'=vy'+v'y you have on the left hand side of vy'+vy=3t^2 vy'+vy In order for vy'+v'y=vy'+vy v' needs to be the same as v This is how we will choose v v=v' assuming y and v are functions of x we have v=dv/dt this equation can be solved by separation of variables

myininaya (myininaya):

dt=dv/v now integrate

myininaya (myininaya):

however if you aren't used to deriving your own integrating factor and you have become reliant upon the form \[v=e^{ \int\limits_{}^{}p(x) dx } \text{ where you have written the differential equation } \\ y'+p(x)y=q(x)\]

myininaya (myininaya):

this is not the answer this just the integrating factor this v is what you need to multiply both sides so that you have (vy)'=vq Then you integrate both sides giving you \[vy=\int vq dx +C \]

OpenStudy (mony01):

what do i do with the intervals?

OpenStudy (mony01):

with this interval [0,1]

myininaya (myininaya):

we will get there after we find y

myininaya (myininaya):

also y(0)=1 is an ordered pair

myininaya (myininaya):

we will use this to find C

myininaya (myininaya):

also you had functions of t not x whatever we will say the variable x represents the variable t

OpenStudy (mony01):

am i doing this good so far y+dy=3t^2 dx

myininaya (myininaya):

Ok I don't think you understand... Let's see if you can following this example: \[\frac{1}{t}y'+y=t^2 \] y is a function of t so first write as \[y'+ty=t^3\] I just multiply t on both sides dt/dt=1 so we already have (ty)'=t^3 integrate both sides with respect to t we have ty=t^4/4+C

myininaya (myininaya):

Here is a harder question \[\frac{1}{t}y'-y=t^2 \] write in the y'+py=q form \[y'-ty =t^3\] multiply both sides by v so that we can sue product rule on left side we get to choose v to be anything (well except 0) \[vy'-vty=vt^3 \] we want to choose this v so we can write the left hand side as vy'+v'y so this means we want -vt=v' so this means we first need to find this v from this equation before we continue we need to know what we multiplied by \[-vt =\frac{dv}{dt} => -t dt = \frac{dv}{v} => -\frac{t^2}{2}=\ln|v| \] So \[v=e^{\frac{-t^2}{2}}\] so we have \[(vy)'=e^{\frac{-t^2}{2}} \cdot t^3\] The integrate both sides you can go ahead and replace the other v as well

myininaya (myininaya):

You can skip all of this though and use the formula but formulas are no fun

myininaya (myininaya):

all of the first order linear differential equations that is y'+py=q can be solved in the same way

myininaya (myininaya):

first find \[v=e^{ \int\limits_{}^{} p(x) dx }\] This is called the integrating factor once you put your equation in this form y'+py=q

myininaya (myininaya):

Then integrate both sides of (vy)'=vq

myininaya (myininaya):

which will lead to \[vy= \int vq dx +C\]

myininaya (myininaya):

but I only recommending using this formula if you actually understand how it works but even then I never like to memorize this formula because it is fun getting there without it

OpenStudy (mony01):

yea im so confused don't i need to use Euler's method?

myininaya (myininaya):

I find this way easiest mostly because I don't remember the way. I would have to go back and do review. I think @rational is watching this like a hawk though. So if he doesn't chime in, I will after I do a quick review.

myininaya (myininaya):

I don't remember the other way*

myininaya (myininaya):

There is also another way where you can guess what y is based on that 3t^2 thing

myininaya (myininaya):

but i need a few minutes or maybe 30 for euler's method

myininaya (myininaya):

so we have f(0,1) <--this represents slope at (0,1)

myininaya (myininaya):

and y'=3t^2-y

myininaya (myininaya):

so y'=3(0)^2-1 so the slope at (0,1) is -1 so f(0,1)=-1

myininaya (myininaya):

The next step is multiply by a step size h hf(y_0)=hf(1)=h(-1) (we will let h=1) so hf(y_0)=-1 where y_0=-1 and t_0=0 does this look familiar so far?

myininaya (myininaya):

then keep evaluating : y_n=y_(n-1)+hf(y_(n-1))

myininaya (myininaya):

but usually i think you use this method to approximation other values

myininaya (myininaya):

Could be wrong about that.

OpenStudy (rational):

hey @mony01 why do you want to use a numeral method when you can solve it analytically ?

myininaya (myininaya):

oops that y_0 is suppose to be 1 (that was a type-0)

OpenStudy (rational):

euler's numerical method gives you a set of points using which you can fit a polynomial/exponential curve.. but i feel you should think of numerical methods only when its not possible to solve it other ways

OpenStudy (rational):

Also, that integrating factor idea was first occurred to Euler only. So please be clear on what you mean by euler's method because Euler is there everywhere. "euler's method" is not a standard phrase. Do you mean Euler's numerical method ?

OpenStudy (mony01):

okay this is the original question for the problem, now im confused on what formula to use. @rational @myininaya Solve the IVP numerically on the suggested interval, if given, using various step sizes. Compare with values of exact solutions when possible.

myininaya (myininaya):

http://www.swarthmore.edu/NatSci/echeeve1/Ref/NumericInt/Euler1.html This is why I think euler's method is used to approximate Are we trying to appoximate a certain value like y(1/2) or something?

OpenStudy (mony01):

I dont really know well the chapter from this problem is all about Euler